Circular No. 7432

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cl IN NGC 3318 Robin Chassagne, Ste. Clotilde, Île de Réunion, reports the discovery of

an apparent supernova (mag 14.8) on two unfiltered CCD images (limiting mag 20) taken with a 0.305-m telescope on May 26.666 and 27.729 UT. SN 2000cl is located at α = 10^{h}37^{m}16^{s}.07, δ = −41^{o}37′47′′.8 (equinox 2000.0), which is 6′′ east and 7′′ south of the center of NGC 3318. Nothing appears at this location on an image obtained by Chassagne on 2000 Feb. 2 (limiting mag 20) or on a U.K. Schmidt (Digital Sky Survey) image taken on 1980 Jan. 24 (limiting mag ∼ 22).

A. Maury, Observatoire de la Côte d’Azur, writes: “L. Vanzi, L. Testi, C. Lidman, and I obtained a fully reduced 30-min spectrum of SN 2000cl with the 1.5-m Danish telescope (+ DFOSC) at the European Southern Observatory on May 27 UT. The spectrum shows the typical features of a type-II supernova: broad Hα, Hβ, and He I at 587.5 nm (the latter with a clear P-Cyg profile). An image and the spectrum may be viewed at

http://sc6.sc.eso.org/~lvanzi/sn3318.html.”

ROSAT ALL-SKY SURVEY FAINT SOURCE CATALOGUE W. Voges, B. Aschenbach, T. Boller, H. Bräuninger, U. Briel, W.

Burkert, K. Dennerl, J. Englhauser, R. Gruber, F. Haberl, G. Hartner, G. Hasinger, E. Pfeffermann, W. Pietsch, P. Predehl, J. Schmitt, J. Trümper, and U. Zimmermann, Max-Planck-Institut für Extraterrestrische Physik, Garching, report: “The ROSAT All-Sky Survey Faint Source Catalogue (RASS-FSC) has been released and is available through the World Wide

Web (at http://wave.xray.mpe.mpg.de/rosat/catalogues/rass-fsc/)

and via anonymous ftp (host ftp.xray.mpe.mpg.de, directory rosat/ catalogues/rass-fsc). This catalogue is derived from the all-sky survey performed during the ROSAT mission in the energy band 0.1–2.4 keV, and 105 924 sources are catalogued, representing the faint extension to the RASS bright-source catalogue (RASS-BSC; cf. IAUC 6420; Voges et al. 1999, A.Ap. 349, 389). The sources contain at least six source photons and have a detection likelihood, −ln (1 − P), of at least 7, where P is the probability of source detection. For each source we provide the ROSAT name, the position in equatorial coordinates, the positional error, the source countrate and error, the background countrate, exposure time, date of observation, two hardness ratios and errors, extent and likelihood of extent, and likelihood of detection. Questions or comments may be directed

to survey@xray.mpe.mpg.de.”

2000 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green